Peter has money in two savings accounts. One rate is 10% and the other is 15%. If he has $950 more in the 15% account and the total interest is $320, how much is in each savings account?
Answer by mananth(16946)

Amount @ 10.00% x
Amount @ 15.00% x + 950
Total interest earned = 320

10.00% x+ 15.00% ( x + 950 )= 320
10.00% x+ 15.00% * x + 142.5 = 320
10.00% x+ 0.15 x + 142.5 = 320
0.10 x+ 0.15 x = 177.5
0.25 x = 177.5
x= 710 Amount @ 10.00%

Amount @ 15.00% = 710 + 950
= 1660